40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 9 00:15:21 2009, Yoichi, Update, PSL, Laser head temperature oscillation laser.png
    Reply  Sat May 9 02:20:11 2009, Jenne, Update, PSL, Laser head temperature oscillation 
       Reply  Sat May 9 15:19:10 2009, rana, Update, PSL, Laser head temperature oscillation Picture_2.png
          Reply  Mon May 11 11:49:20 2009, steve, Update, PSL, MOPA cooling water lines are backwards htempvar80d.jpg
             Reply  Thu May 21 18:07:32 2009, rana, Update, PSL, Screw on Needle valve loosened Untitled.png
Message ID: 1573     Entry time: Mon May 11 11:49:20 2009     In reply to: 1570     Reply to this: 1617
Author: steve 
Type: Update 
Category: PSL 
Subject: MOPA cooling water lines are backwards 

Quote:
This is 8 days of 10-minute trend.

DTEC is just the feedback control signal required to keep the NPRO's pump diode at a constant temperature.
Its not the amplifier or the actual NPRO crystal's temperature readout.

There is no TEC for the amplifier. It looks like to me that by opening up the flow to the NPRO some more
we have reduced the flow to the amplifier (which is the one that needs it) and created these temperature
fluctuations.

What we need to do is choke down the needle valve and ream out the NPRO block.




I have measured the "input" line temp at the MOPA box 10 C and the "out" line 8 C

This must be corrected.

However look at the 80 days plot of operation where the head temp variation is nothing new
Attachment 1: htempvar80d.jpg  80 kB  | Hide | Hide all
htempvar80d.jpg
ELOG V3.1.3-